to find the height of a tower, a surveyor positions a transit that is 2 meters tall at a spot 40 meters from the base of the tower. She measures the angle of elevation to the top of the tower to be 46 degrees. What is the height of the tower, to the nearest meter?
Answer by mananth(16946)

to find the height of a tower, a surveyor positions a transit that is 2 meters tall at a spot 40 meters from the base of the tower. She measures the angle of elevation to the top of the tower to be 46 degrees. What is the height of the tower, to the nearest meter?
...
Let the height of the tower be x m
distance from base = 40 m
the angle of elevation from 2 m height = 46 deg.
so the vertical distance measured will be x-2
the vertical height will be x-2 m
...
Tantheta = vertical /base
Tan (46)=(x-2)/40
40*tan(46)=x-2
40*tan(46)+2 = x
x = 43 meters
...
m.ananth@hotmail.ca